### v1

# RadOnc-GPT: An Autonomous LLM Agent for Real-Time Patient Outcomes Labeling at Scale

Abstract
Purpose: Manual labeling limits the scale, accuracy, and timeliness of patient outcomes research in radiation oncology. We present RadOnc-GPT, an autonomous large language model (LLM)–based agent capable of independently retrieving patient-specific information, iteratively assessing evidence, and returning structured outcomes. Our evaluation explicitly validates RadOnc-GPT across two clearly defined tiers of increasing complexity: (1) a structured quality assurance (QA) tier, assessing the accurate retrieval of demographic and radiotherapy treatment plan details, followed by (2) a complex clinical outcomes labeling tier involving determination of mandibular osteoradionecrosis (ORN) in head-and-neck (H&N) cancer patients and detection of cancer recurrence in independent prostate and head-and-neck cancer cohorts requiring combined interpretation of structured and unstructured patient data. The QA tier establishes foundational trust in structured-data retrieval, a critical prerequisite for successful complex clinical outcome labeling.
Methods: A wrapper program called LLM Task Streaming sequentially supplied patient IDs and specific task instructions to RadOnc-GPT. Initially, RadOnc-GPT completed the structured QA task using demographic and radiotherapy treatment plan course data exclusively. In the second tier, RadOnc-GPT autonomously combined structured electrical health record (EHR) data with unstructured clinical notes, radiology, and pathology reports to determine clinical outcomes. Ground-truth labels for ORN and cancer recurrence were retrospectively generated by expert radiation oncologists. Discrepancies between RadOnc-GPT’s outputs and ground-truth labels underwent independent adjudication by independent radiation oncologists, classifying each discrepancy into one of three categories: model error, ground-truth error, or indeterminate.
Results: Across 895 unique patients, RadOnc-GPT matched all six demographic fields for all 500 patients in the QA tier (100%) and accurately reproduced radiation-course counts in 497/500 (99.4%), demonstrating robust retrieval fidelity. In the complex patient outcomes labeling tier, initial accuracy improved significantly post-adjudication: for ORN determination (233 patients) accuracy rose from 84.5% to 95.2%, prostate cancer recurrence detection (80 patients) improved from 92.5% to 95.0%, and H&N recurrence detection (82 patients) improved from 92.7% to 96.3%. Among 48 initial discrepancies, adjudication revealed 30 (63%) previously unrecognized ground-truth errors, 13 genuine model errors, and 5 indeterminate cases.
Conclusions: RadOnc-GPT reliably retrieves foundational structured data and subsequently generalizes complex clinical outcome labeling tasks effectively, notably using a single cancer recurrence detection prompt across multiple disease sites. Its high recall performance minimizes clinically critical false negatives, while its dual capacity as labeler and auditor significantly enhances registry data integrity. Thus, RadOnc-GPT enables scalable, trustworthy, and real-time curation of radiation-oncology research datasets.
1 Introduction
Large-language models (LLMs) have progressed rapidly from general-purpose chatbots to highly specialized systems capable of matching or surpassing human experts across multiple clinical tasks[1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Specifically, in radiation oncology GPT-4 has attained expert-level performance on radiation-oncology physics exams [11, 12], broad clinical knowledge assessments [13, 14, 15], clinical task management such as in-basket response [16], automated drafting of payer-appeal documents [17], and harmonizing naming conventions of anatomical structures [18]. In addition, recently open source LLMs have been fine-tuned to improve the performance of important clinical tasks in radiation oncology [19]. Further improvements have been demonstrated through retrieval-augmented generation (RAG) approaches, which significantly enhance accuracy in clinical trial screening applications [20, 21, 22, 23, 24].
Parallel studies in oncology-specific natural language processing (NLP) have demonstrated the successful extraction of detailed clinical information such as pathology findings [25], computed tomography (CT) imaging phenotypes [26], key entities from discharge letters [27], toxicity assessments [28], patient comorbidities [29], outcomes from lung-ablation procedures [30], and patient-reported outcomes [31]. Recent multi-modal models that integrate imaging and textual data have further expanded capabilities to tasks like tumor contouring [32] and survival prediction [33]. However, a significant limitation of most existing NLP approaches and registry curation efforts is their inherent assumption that institutional registry labels represent reliable gold standards. Recent audits, such as a multi-year review of an academic head-and-neck cancer registry, suggest otherwise, revealing significant latent errors, with up to half of true recurrence events potentially missing from automated flags [34]. Correcting these errors via traditional manual review would require substantial staffing resources, highlighting the urgent need for more efficient, scalable methods for auditing and curating clinical outcomes data.
To address these critical limitations, we developed RadOnc-GPT, a GPT-4o–based autonomous agent capable of retrieving structured and unstructured clinical data directly from institutional databases, iteratively assessing the evidence, and synthesizing structured clinical outcomes. In order to evaluate RadOnc-GPT, we designed a two-tier strategy differing by their level of complexity: first, a relatively simple structured data retrieval task aimed at establishing the agent’s structured information retrieval capability. Because this task involves reproducing structured data, manual review is not required and therefore makes this evaluation step a good candidate for quality assurance purposes. After demonstrating reliable structured-data reproduction, we increase the complexity of the tasks, evaluating RadOnc-GPT’s ability to mix structured and unstructured patient data towards performing the task. These high-complexity tasks included determination of mandibular osteoradionecrosis (ORN) in head-and-neck cancer patients and cancer recurrence detection in separate prostate and head-and-neck cancer patient cohorts. A noteworthy design feature of our evaluation was using an identical cancer recurrence detection method (identical input prompt) across two disease sites—prostate cancer and head-and-neck cancer—testing RadOnc-GPT’s ability to generalize clinical reasoning across disease sites.
Through this staged evaluation approach, we aim to answer three key questions:
  1. 1.
    Can an autonomous LLM-based agent reliably retrieve and interpret structured clinical data from institutional databases?
  2. 2.
    Once baseline structured retrieval accuracy is confirmed, can the agent successfully perform complex clinical outcomes labeling using both structured and unstructured data?
  3. 3.
    Can such an agent simultaneously identify latent errors in existing institutional registry labels, thus serving a dual function of patient outcome labeling and real-time data auditing?
Answering these questions clearly and rigorously establishes autonomous LLM agents as a potentially transformative solution for near-real-time, scalable and accurate clinical outcomes labeling.
2 Methods
2.1 RadOnc-GPT System Design
Figure 1 illustrates the architecture of RadOnc-GPT, which integrates both internal and external data resources. Internal connections include Mayo Clinic’s radiation oncology–specific database system (Minnesota, Arizona, Florida), Aria (Varian Medical Systems, Palo Alto, CA), and enterprise electronic health record (EHR) database systems interfacing with Epic (Epic Systems, Verona, WI). RadOnc-GPT further connects with public data resources including PubMed, ClinicalTrials.gov, and the National Cancer Institute (NCI) Common Terminology Criteria for Adverse Events (CTCAE), each connected via their public APIs. RadOnc-GPT is designed as an LLM-based agent, defined here as a system capable of conducting multi-turn conversations while autonomously determining which functions to call and when to stop responding. The RadOnc-GPT system makes use of the many data sources by calling upon whitelisted functions that retrieve, process, and return data to GPT-4o (OpenAI) or Gemini 2.5 Pro (Google). The specifications of these functions—including their inputs and outputs—are explicitly detailed within the system prompt so as to promote good decision making on the part of the LLM in deciding which functions to run and when to continue or stop responding.
Figure 1: RadOnc-GPT Agent Architecture.
2.2 Data Retrieval Functions
Table 1 summarizes the whitelisted functions available to RadOnc-GPT. When invoked, each function executes and returns outputs directly to the model. These outputs are designed to provide a coherent, structured representation of the underlying data, thereby maximizing the LLM’s ability to interpret and reason over the information. If no relevant data were found by a particular data retrieval function, the function would explicitly indicate that no data could be found for the given inputs and patient.
Importantly, RadOnc-GPT does not employ conventional retrieval-augmented generation (RAG). In RAG, embedded documents are matched to queries via vector similarity and delivered back to the model as text chunks—a strategy developed to surface relevant passages from vast, poorly organized unstructured corpora such as the internet, which often exceed model context limits. By contrast, RadOnc-GPT leverages the fact that patient data within Epic is unstructured yet systematically organized and indexed. Clinical notes, for example, are timestamped, provider-signed, and tagged with metadata such as medical specialty and treating department, making them straightforward to retrieve without generic RAG pipelines. As LLMs continue to support dramatically larger contexts (e.g., GPT-4o accommodates up to 1 million tokens) while inference costs per token decline (\$2.50 per 1 million input tokens for OpenAI’s standard GPT-4o now), the necessity of general-purpose RAG approaches, which introduce unnecessary complications to the system, further diminishes in settings like RadOnc-GPT, where targeted, well-structured data retrieval suffices. For instance, when managing a prostate cancer patient, RadOnc-GPT retrieves all radiation oncology and urology clinical notes relevant to that patient.
To avoid overwhelming the model with unnecessarily lengthy inputs, RadOnc-GPT employs a large set of highly specific retrieval functions. The guiding principle is that greater function granularity reduces the volume of data returned on a per-function basis. In rare cases where retrieved content exceeds the maximum allowable context length, a controlled pruning strategy is applied. The system prompt remains intact, while prior messages are selectively truncated. Because all retrieval functions return results in reverse chronological order (newest first, oldest last), pruning proceeds by removing fixed-size batches of words from the ends of the oldest messages. After each pass, the token count is recalculated; if the context remains too long, pruning continues until the message reaches a minimum threshold, at which point the process shifts to the next oldest message. Messages already at or below a skip threshold are preserved until all messages fall below that limit; if the conversation still exceeds the token budget, the threshold is lowered stepwise and pruning resumes. For this study, parameters were set as follows: maximum conversation history = 95,000 tokens, words removed per pass = 250, minimum threshold = 10,000 tokens, and threshold decrement = 2,500 tokens.
Table 1: RadOnc-GPT whitelisted function categories.
Category Purpose / Data Retrieved Representative Functions
Patient-oriented data retrieval Demographics, treatments, diagnoses, notes, reports, messages, appointments get_patient_details, get_patient_treatment_details, get_patient_diagnosis_details, get_patient_clinical_notes, get_patient_radiology_reports, get_patient_pathology_reports, get_patient_inbasket_messages, get_patient_appointments, get_physician_appointments
Clinical trials Identify recruiting trials and fetch eligibility criteria from clinicaltrials.gov get_list_of_clinical_trials, get_eligibility_criteria
Statistical information Generate cohort-level population statistics / search URLs get_patient_population
PSTAR (proton range) Compute proton stopping power, CSDA and projected ranges get_pstar_data, get_pstar_material_list
PubMed Search biomedical literature, retrieve summaries and abstracts pubmed_search, pubmed_summary, pubmed_fetch
DICOM Send DICOMs, monitor active streams, clear logs send_dicoms_to_server, get_dicom_streams_info, clear_dicom_streams_logs
CTCAE (AE grading) List radiation-therapy adverse-event codes and grading details get_radiation_therapy_ae_codes, get_ctcae_details_by_ae_code
Basic math Perform elementary arithmetic inside the agent loop add, subtract, multiply, divide
2.3 Two-Tier Evaluation Scaffold
RadOnc-GPT’s capability to extract patient outcomes was evaluated in two distinct tiers of escalating complexity (see Figure 2). This two-tiered approach establishes foundational accuracy in basic retrieval tasks first, followed by more clinically relevant tasks of clinical outcomes labeling:
  • Tier 1 (Data extraction from structured data): In the first evaluation, RadOnc-GPT was tasked with retrieving and reproducing structured data elements, including patient ID, name, sex, race, ethnicity, and treatment planning details such as course IDs, International Classification of Diseases (ICD) codes, plan IDs, and radiation types for each course. RadOnc-GPT’s outputs were directly compared with corresponding values extracted from the underlying database. Because this task relied exclusively on structured fields, no adjudication process was required, as the database entries were considered the ground truth.
  • Tier 2 (Patient outcomes labeling from structured and unstructured patient data): In the second evaluation, RadOnc-GPT was assessed on three clinically complex outcome-labeling tasks: (1) detection of mandibular osteoradionecrosis (ORN) in patients with head-and-neck cancer following radiation therapy, (2) detection of cancer recurrence in prostate cancer patients after radiation therapy, and (3) detection of recurrence in head-and-neck cancer patients after surgery. These tasks required RadOnc-GPT to autonomously synthesize both structured data (e.g., demographics, treatment planning details, diagnosis codes) and unstructured data (e.g., clinical notes, radiology reports, pathology reports). For recurrence detection, the same prompt design was applied across prostate and head-and-neck cohorts, allowing evaluation of RadOnc-GPT’s ability to generalize across disease sites and treatment modalities. The outcomes labels generated by RadOnc-GPT were compared with ground truth labels that were acquired through careful manual review by experienced physicians. Our study considers that ground truth may not be accurate. Each discrepancy between RadOnc-GPT and the ground truth was therefore adjudicated by an independent physician for a final result.
Figure 2: Tiered difficulty of tasks. Two-Tier evaluation scaffold where tier 1 is relatively simple, structured data-only extraction task, and tier 2 is a difficult task requiring structured and unstructured data. Tier 1 evaluation does not require manual review; Tier 2 does.
2.4 LLM Task Orchestration
To enable large-scale evaluation, we developed an external orchestration framework called LLM Task Streaming, designed to allow for processing a population of patients with RadOnc-GPT. See Figure 3. This system sequentially supplied RadOnc-GPT with patient identifiers from each cohort along with task-specific instructions (prompts). The model’s outputs were returned as structured JSON objects, which were automatically parsed and consolidated into cohort-level CSV files for downstream analysis and physician adjudication. The framework supported parallel processing, handling up to four patients concurrently. In practice, overall throughput was constrained primarily by the inference speed of GPT-4o, with each patient’s data typically processed within 10–20 seconds depending on the volume of retrieved records.
Figure 3: LLM Task Streaming Orchestration. External controller (LLM Task Streaming) sequentially supplying patient IDs and task instructions (prompts) to RadOnc-GPT, aggregating structured JSON objects.
2.5 Prompts
To execute a given task, RadOnc-GPT requires two inputs: the patient identifier and a task-specific prompt. The full set of prompts used for each task is provided in the Appendix. In general, all prompts followed a consistent design strategy: they specified the overall task, outlined which patient data should be retrieved, and included task-specific instructions. These task-specific components could incorporate definitions, clinical criteria, or reasoning strategies (e.g., chain-of-thought guidance) to ensure accurate and consistent outputs. The process of developing the prompts involved frequent observation and feedback from experienced physicians.
2.6 Cohorts and Ground Truth Labeling
Four non-overlapping cohorts were selected for evaluation (total $n=895$). Table 2 summarizes the number of patients (total, positive, and negative) included in the Tier 2 (complex) tasks. For the Tier 1 task, the cohort consisted of $n=500$ patients randomly selected from radiation therapy patients treated between 2015 and 2025. The ground truth for this evaluation was derived directly from the institutional Aria database.
Table 2: Patient cohort summary and baseline ground-truth label counts (pre-adjudication).
Cohort Total $n$ Positive (n) Negative (n)
ORN detection (head-and-neck) 233 34 199
Prostate recurrence detection 80 40 40
Head-and-neck recurrence detection 82 48 34
Counts reflect baseline labels prior to adjudication; post-adjudication revisions are reported in the Results section.
Mandibular osteoradionecrosis (ORN) cohort.
Selection: Head-and-neck cancer patients were included if they met prespecified criteria: adults (>>19 years) treated between April 2013 and August 2019 with fraction sizes of 1.2–2.2 Gy and a total dose >>30 Gy, who subsequently developed mandibular osteoradionecrosis (ORN) after their initial radiotherapy course. Patients with recurrent cancer were excluded.
Ground truth: ORN status was established using CTCAE v4 osteonecrosis-of-the-jaw criteria and graded according to the Marx system. Positivity required documentation of ORN within the irradiated volume, supported by clinical notes, imaging, and lesion photographs. Alternative etiologies (e.g., infection, surgery, dental extraction, lymphoma, or disease outside the radiation field) were excluded.
Head-and-neck cancer recurrence cohort.
Selection: Head-and-neck cancer patients with prior radiation therapy.
Ground truth: Recurrence status was determined through chart review by an experienced clinical data abstractor (¿$15$ years), using predefined criteria established by a surgical oncologist. Ambiguities were resolved through weekly review sessions with the oncologist. All head-and-neck primary sites and histologies were eligible for inclusion.
Prostate cancer recurrence cohort.
Selection: Prostate cancer patients treated with definitive, curative-intent radiotherapy.
Ground truth: Recurrence was defined as any documented post-treatment progression of prostate cancer and was labeled if either criterion was met: (1) biochemical recurrence (PSA nadir + 2 ng/mL) or a clinician-determined, intervention-triggering PSA rise; and/or (2) clinical recurrence—local, regional, distant nodal, or metastatic progression confirmed by imaging or biopsy. Only prostate cancer recurrences were considered when other primary cancers were present.
2.7 Adjudication Procedure
For the tier 2 (complex) outcomes tasks, discrepancies between RadOnc-GPT’s results and baseline labels were independently adjudicated by oncologists with domain expertise specific to each clinical scenario (prostate cancer, head-and-neck cancer, mandibular osteoradionecrosis), one oncologist per dataset. Oncologists classified discrepancies explicitly into three categories:
  1. 1.
    RadOnc-GPT Correct: Baseline label was determined erroneous (ground-truth error).
  2. 2.
    Baseline Label Correct: RadOnc-GPT’s decision was erroneous (model error).
  3. 3.
    Indeterminate: Insufficient evidence was available to reach a clear adjudication; these cases were excluded from final metrics.
Accuracy, precision, recall, and F1-score metrics were computed before and after the adjudication process.
3 Results
3.1 Tier 1 Task - Structured Data Extraction (QA)
In the initial structured extraction QA tier, RadOnc-GPT demonstrated exceptionally high fidelity in accurately retrieving structured demographic and treatment-related information. Across all 500 evaluated patients (totaling 3,000 individual demographic fields), RadOnc-GPT matched the demographic database records perfectly (100%). Additionally, RadOnc-GPT accurately reproduced the treatment planning information in 497 out of 500 cases (99.4%). These results established foundational trust in RadOnc-GPT’s basic retrieval capability, crucial for subsequent complex clinical outcomes labeling based on structured and unstructured patient data.
3.2 Tier 2 Tasks - Complex Clinical Outcomes Labeling based on Structured and Unstructured patient data
Mandibular Osteoradionecrosis (ORN) in Head-and-Neck Patients:
Among 233 head-and-neck patients evaluated for ORN, RadOnc-GPT initially achieved an accuracy of 84.5%, with high recall (88.2%) but lower precision (48.4%). After adjudication by an independent radiation oncologist, overall accuracy significantly improved to 95.2%. The adjudication process revealed that of the 36 initial discrepancies (32 false positives, 4 false negatives):
  • 25 cases (69%) were previously unrecognized ground-truth errors, primarily due to outdated or incorrectly documented clinical information.
  • 7 discrepancies were confirmed as genuine model errors.
  • 4 cases were indeterminate due to insufficient evidence, and were excluded from final analysis.
Post-adjudication, RadOnc-GPT achieved 100% recall, eliminating all false negatives.
Prostate Cancer Recurrence:
In the prostate cancer recurrence cohort (80 patients), RadOnc-GPT initially reached an accuracy of 92.5%, with precision and recall also both at 92.5%. Post-adjudication, the accuracy improved to 95.0%. Of the 6 initial discrepancies (3 false positives, 3 false negatives):
  • 2 cases (33%) were found to be ground-truth errors, both involving small-volume lymph node recurrences that were initially missed during manual chart review.
  • 4 discrepancies were confirmed as genuine model errors.
  • No indeterminate cases were identified in this cohort.
The post-adjudication recall remained at 92.5%, demonstrating reliable clinical utility for identifying prostate cancer recurrence.
Head-and-Neck Cancer Recurrence:
For the head-and-neck cancer recurrence cohort (82 patients), initial accuracy was 92.7%, with a precision of 92.0% and recall of 95.8%. After independent adjudication, accuracy significantly increased to 96.3%. The 6 initial discrepancies (4 false positives, 2 false negatives) were adjudicated as follows:
  • 3 cases (50%) were determined to be ground-truth errors, reflecting incorrect or outdated documentation.
  • 2 cases were genuine model errors.
  • 1 case was indeterminate and thus excluded from the final analysis.
Final recall was excellent at 97.9%, reflecting a clinically important reduction in false negatives.
3.3 Summary of Adjudication Results Across Complex Patient Outcomes Labeling Tasks based on Structured and Unstructured Patient Data
Across three complex patient outcomes labeling tasks (ORN in head–and–neck, prostate cancer recurrence, head–and–neck cancer recurrence; total $n=395$ before adjudication), expert review of all discordant cases ($n=48$) found 30 (63%) baseline labeling errors, 13 model errors, and 5 indeterminate cases. Post–adjudication, false negatives fell from 9 to 4, and pooled (micro–averaged) performance improved from Precision/Recall/F1/Accuracy of 74.3/92.6/82.5/87.8% to 90.3/97.0/93.6/95.4% (Table 3). Note that post–adjudication denominators exclude indeterminate cases for the affected patient cohorts.
Table 3: Confusion counts and performance metrics before and after adjudication. Totals are pooled (micro–averaged) across tasks. Post–adjudication NN excludes indeterminate cases (ORN: 4; H&N recurrence: 1).
Task Adj. NN TP FP FN TN Pr.(%) Re.(%) F1(%) Ac.(%)
ORN (H&N) Before 233 30 32 4 167 48.4 88.2 62.5 84.5
After 229 48 11 0 170 81.4 100.0 89.7 95.2
Rec. (Prost.) Before 80 37 3 3 37 92.5 92.5 92.5 92.5
After 80 37 1 3 39 97.4 92.5 94.9 95.0
Rec. (H&N) Before 82 46 4 2 30 92.0 95.8 93.9 92.7
After 81 46 2 1 32 95.8 97.9 96.8 96.3
Total Before 395 113 39 9 234 74.3 92.6 82.5 87.8
After 390 131 14 4 241 90.3 97.0 93.6 95.4
4 Discussion
This study investigated whether an autonomous, self-retrieving LLM agent can simultaneously curate patient outcomes and audit patient data quality within a realistic clinical setting. We adopted a two-tier evaluation strategy by design: first establishing fidelity in structured data retrieval before advancing to the more complex task of patient outcome labeling. In Tier 1, RadOnc-GPT successfully reproduced structured data alone, achieving exact matches for all demographic fields and accurately reproducing per-course treatment information in 497 of 500 patients, including course identifiers, ICD codes, delivered plan identifiers, and radiation modality. Because these comparisons were made against structured database values from Aria, no manual review was required, making this approach well-suited for routine quality assurance. Importantly, this QA test also surfaced a systematic error early in our evaluation: course IDs with leading numbers were occasionally truncated by the LLM. Identifying this issue allowed us to implement corrections and strengthen the retrieval pathway. Thus, Tier 1 not only served as a validity checkpoint but also provided a mechanism for hardening the retrieval process before progressing to more complex evaluative tasks.
For Tier 2 tasks, after adjudication, recall reached 100% for ORN detection, 92.5% for prostate cancer recurrence detection, and 97.9% for head-and-neck cancer recurrence detection. In the context of surveillance and registry curation, false negatives are typically unobserved whereas false positives remain reviewable; thus, a recall-oriented operating point is preferable. The ORN finding—absence of false negatives—is particularly noteworthy, as missed late toxicities carry substantial clinical consequences. Two design choices likely contributed to this performance: (1) prompts explicitly requiring evidence both for and against each Marx stage, and (2) function-call–guided retrieval with task-specific temporal bounds, which maintains clinical context while filtering out irrelevant text. Together, these elements mitigate risks such as fragmentation of clinically linked facts across retrieval chunks and the limitations of generic embeddings, which often underperform in radiotherapy-specific domains.
Adjudication reframed model-versus-label disagreements. Of 48 initial discrepancies, 30 (63%) were attributable to baseline labeling errors, 13 to model-generated errors, and 5 remained indeterminate. After excluding indeterminate cases, the adjusted accuracies of the original institutional labels were 95.2% for ORN, 97.5% for prostate cancer recurrence, and 96.3% for head-and-neck cancer recurrence. Sutton et al. reported for head-and-neck cancer a registry accuracy of 89.4% with sensitivity 61% and specificity 99%, further noting that longer time-to-recurrence was associated with false negatives (median 1,007 vs. 414 days for true positives) [34]. Against this backdrop, RadOnc-GPT achieved higher recall while maintaining high accuracy, and it surfaced labeling errors that would otherwise have persisted. Although the comparison is indirect—owing to differences in definitions and cohorts—the findings suggest that autonomous agents can enhance completeness relative to conventional registry processes, while simultaneously generating an audit trail to support adjudication.
Throughput supports near-real-time use. With four-way parallelism and per-patient latencies of 10-20 seconds, most cohorts can be processed within a single day. Outputs are returned as structured JSON, enabling seamless integration into dashboards, analytics pipelines, and trial-matching platforms. This operational profile is important because registry accuracy is time-sensitive: delays in documentation contribute to missed events in conventional workflows [34]. A nightly or on-demand run can substantially shorten this gap. Notably, the same cancer recurrence detection prompt was applied to both prostate and head-and-neck cancer, suggesting that disease-site generalization is feasible with foundational LLM-based systems.
Limitations of this study include its reliance on single-center data and single-reviewer adjudication for each dataset; multi-reviewer panels could shift borderline determinations. The underlying model is proprietary (GPT-4o), and both governance considerations and financial cost may constrain adoption, with continuous institution-wide runs remaining cost-sensitive at present. These factors motivate future benchmarking against open-source models, exploration of incremental recomputation and caching strategies, and deployment patterns that respect privacy requirements (including federated approaches). Finally, although the agent generated rationale-rich outputs that facilitated adjudication, we did not evaluate human–AI conflict resolution policies in live operations; such mechanisms will be essential for routine deployment.
5 Conclusions
RadOnc-GPT demonstrates that a self-retrieving GPT-4o agent can:
  1. 1.
    flawlessly reproduce fundamental structured data, establishing trust in its retrieval chain;
  2. 2.
    label complex late-toxicity and cancer recurrence endpoints with near-expert performance;
  3. 3.
    simultaneously surface hidden errors in institutional databases, creating a virtuous cycle of model-assisted quality improvement.
Because the approach is prompt-driven, requires no model fine-tuning, and leverages granular yet generic API hooks, it is readily extensible to other oncology centers and disease sites. We envision autonomous LLM agents functioning as continuous auditors—running nightly against new patients, updating outcomes registries in real time, and allowing clinicians to focus on judgment rather than data wrangling. The anticipated result is richer and timelier evidence to inform precision radiotherapy, ultimately advancing patient care.
Acknowledgements
Institutional policy required that any physician with greater than 21 patients involved be included as an author.
Disclosures
Appendix A Prompts
The following prompt was used for the structured QA task:
Retrieve the patient details and patient treatment details for patient
{patient_id}. Based on the delivered treatment details, report the following
information, like this:
    "patient_id": "12345678",
    "first_name": "first name",
    "last_name": "last name (surname)",
    "sex": "male"/"female",
    "race": "put race here",
    "ethnicity": "put ethnicity here",
    "delivered_courses": [
            "course_id": "course ID 1",
            "icd_codes": ["C61", "C75.1", ...],
            "delivered_plan_ids": ["plan ID 1", "plan ID 2", ...],
            "radiation_type": "proton" or "photon" or "electron"
            "course_id": "course ID 2",
            "icd_codes": ["C61", "C75.1", ...],
            "delivered_plan_ids": ["plan ID 1", "plan ID 2", ...],
            "radiation_type": "proton" or "photon" or "electron"
The following prompt was used for determining whether the patient experienced ORN:
Determine whether patient {patient_id} has ever experienced
osteoradionecrosis (ORN). If the patient did experience ORN,
then grade the severity (using the Marx staging system).
Osteoradionecrosis (ORN) of the mandible refers to a condition
occurring within the radiation treatment (RT) field in the head
and neck region, where mucosal breakdown or delayed healing
leads to persistent exposure of the mandibular bone. Key criteria
include bone exposure lasting for at least three months, evidence
of necrotic (dead) bone, and the absence of recurrent tumor or
metastases at the affected site. Additionally, ORN may present
with radiographic evidence of bone necrosis even when the
overlying mucosa remains intact. In order to determine whether
the patient experienced ORN, first retrieve the patient details,
patient diagnosis details, patient treatment details, radiology
reports, pathology reports, and patient clinical notes (note_type
of radiology, pathology, surgery, radiation_oncology, and ENT).
Provide a detailed summary of the retrieved data, in the context
of determining whether the patient experienced ORN, in a section
called Patient History.
Next, in a section called Data Quantity, quantify the total
amount of patient data that was retrieved by summing the ’number
of records count’ from each of the retrieved patient data (not
including patient details, patient diagnoses details, or patient
treatment details records).
Next, grade the severity of ORN in a section called Marx Staging.
The Marx staging system for grading ORN, proposed by Robert E.
Marx in an article entitled "Osteoradionecrosis: A New Concept
of Its Pathophysiology", is as follows: "Stage 1 is defined as
exposed alveolar or mandibular bone without pathologic fracture.
If hyperbaric oxygen therapy (HBO) is used and the exposed bone
responds positively as a result, this is an indication of stage
1; Stage 2 is disease which does not respond to HBO therapy if
it is given, and requires sequestrectomy and saucerization; and
Stage 3 is full-thickness bone damage or pathological fracture,
usually requiring complete resection and reconstruction with free
tissue." List any evidence for/against each stage (1, 2, or 3).
If none of the criteria is met for any stage, as defined, then we
shall define this as stage 0 (no evidence of ORN).
Finally, in a section called Concluding Remarks, make the case
for or against the patient having ORN. After your concluding
remarks, **provide the results using the following format**:
    ’stage’: ’0, 1, 2, or 3’,
    ’total number of records’: ’10, 14, 25, 115, etc.’
The following prompt was used for determining whether the patient had cancer recurrence:
#### Task
I need your help in determining whether patient {patient_id} has
had cancer recurrence at any point in their history relating to
one of their prior cancer diagnoses.
#### Definition of Cancer Recurrence
According to the National Cancer Institute, recurrent cancer is
cancer that has recurred (come back), usually after a period of
time during which the cancer could not be detected. The cancer
may come back to the same place as the original (primary) tumor
or to another place in the body, but is only considered
recurrence if it is related to the primary tumor. If it is a
new tumor that is unrelated (a new, secondary primary tumor),
then it is not recurrence relating to the original primary
tumor. However, there is also the possibility that recurrence
occurred relating to the secondary primary tumor.
#### Data Retrieval
To make a determination of whether the patient experienced cancer
recurrence in their history, you’ll need to retrieve the following
data...
    * Patient treatment details
    * Radiation oncology notes (‘note_type‘ = ‘radiation_oncology‘)
    * Pathology notes (‘note_type‘ = ‘pathology‘)
    * Radiology notes (‘note_type‘ = ‘radiology‘)
    * Urology notes (‘note_type‘ = ‘urology‘... wait until you see
      diagnosis and only retrieve if the diagnosis is related to
      urology)
    * ENT notes (‘note_type‘ = ‘ent‘... wait until you see diagnosis
      and only retrieve if the diagnosis is related to head and neck
      cancers)
    * Radiology reports
When pulling any of the radiation oncology notes, the pathology
notes, radiology notes, urology notes, ENT notes, and the radiology
reports, use the ‘date_minimum‘ input (I added ‘date_minimum‘ as a
new input for these data retrieval functions, format is %Y-%m-%d
%H:%M:%S) and retrieve them all simultaneously using the last
treatment date of the earliest treatment course as the
‘date_minimum‘. Because you need the treatment details and the
diagnoses details before retrieving certain patient data, you should
retrieve the treatment details and diagnoses details first. Go ahead
and retrieve them. Next, remark as to when the last treatment date
of the first treatment course occurred, what the diagnoses are, and
whether the urology or ENT clinical notes are needed (based on
whether the diagnoses are urology-related or head-and-neck-related
- note that prostate cancer is related to urology). After that,
depending on the last treatment data and diagnosis, retrieve the
rest of the patient data. If the patient does not have treatment
details, then do not use the ‘date_minimum‘ input.
#### Instructions for Determining Cancer Recurrence
Before providing a yes/no answer on whether the patient had cancer
recurrence...
    * Summarize each dataset that was retrieved (including the
      treatment details) in a section called "Summary of Retrieved
      Data". Create a subsection for each summary.
    * Summarize the diagnoses that were treated noting the treatment
      dates in a section called "Summary of Diagnoses".
    * Construct a timeline of important events that include the
      diagnostic imaging reports, pathology imaging/lab reports, the
      actual diagnosis of the disease(s) with onset date(s), the
      treatment of the disease(s), follow-up notes (note which
      diagnosis/treatment the follow-up is referring to (most likely
      the most recent one)) and any other important information
      relating to the progression of disease over time or lack
      thereof in a section called "Disease Timeline". You must
      include PSA measurements in the timeline if the patient had
      prostate cancer.
    * Next, we want to consider an argument that supports recurrence
      and an argument that does not support recurrence in a section
      called "Argument For Recurrence/Argument Against Recurrence".
      First, write a paragraph (starting with "The most plausible
      argument supporting the case that cancer recurrence occurred in
      their history is...") that describes the most plausible
      explanation for how the cancer recurred in their history,
      progressing from one diagnosis to another (consider spatial
      proximity and time proximity). Next, write a paragraph (starting
      with "The most plausible argument supporting the case that
      cancer recurrence did not occur in their history is...") that
      describes the most plausible explanation for how/why the cancer
      did not recur in their history. You must write both of these
      paragraphs unless there was only one diagnosis in their history
      (in which case recurrence was very unlikely).
    * Finally, explain your reasoning as to which choice is most
      plausible, weighing the evidence, followed by a clear statement
      of your final decision in a section called "Concluding Remarks".
    * Provide the answer in an "Answer" section using this JSON format:
‘‘‘json
    "recurrence": "yes/no"